OpenStudy (anonymous):

Check my work. Please!! The formula for determining interest compounded monthly is A = P(1 +r/12)^12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,000 at an interest rate of 10% for 3 years, while Jenny invests $1,000 at an interest rate of 5% for 6 years. Determine the amount of return gained by Jimmy and Jenny. In complete sentences, summarize your results.

OpenStudy (anonymous):

Jimmy: A=$1,000(1+.1/12 )^36 =$1,346.18 Jenny: A = $1,000(1 +.05/12)^7=$4,413.15 Jimmy gains: $1,346.18-1,000 =$346.18 Jenny gains: $4,413.15-1,000=$3,413.15

OpenStudy (anonymous):

Yay! So is my work right?!

OpenStudy (anonymous):

hjmmmmmm

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

some of the answers are off for sure

OpenStudy (anonymous):

Well can you help me fix them?? I'm so confused.

OpenStudy (anonymous):

Jenny: A = $1,000(1 +.05/12)^7=$4,413.15 is wrong

OpenStudy (anonymous):

10% is 0.1

OpenStudy (anonymous):

Yes then I subtracted 1,000 for the answer?

OpenStudy (anonymous):

the exponent should be \(30\)

OpenStudy (anonymous):

why

OpenStudy (anonymous):

and the calculation was also wrong, you must have entered it in the calculator incorrectly

OpenStudy (anonymous):

you can't invest $1000 at 5% for 6 years and make $4000!

OpenStudy (anonymous):

should be \[A=1000\times (1+\frac{.05}{12})^{12\times 6}\]

OpenStudy (anonymous):

or \[A=1000\left(1+\frac{.05}{12}\right)^{72}\]

OpenStudy (anonymous):

its about $1349 http://www.wolframalpha.com/input/?i=1000%281%2B.05%2F12%29^72

OpenStudy (anonymous):

that was jenny

OpenStudy (anonymous):

Jenny**

OpenStudy (anonymous):

first one looks right, i will check it

OpenStudy (anonymous):

i am getting $1348 for the first one, but it is close http://www.wolframalpha.com/input/?i=1%2C000%281%2B.1%2F12+%29^36+

OpenStudy (anonymous):

i think you had the right formula and just entered it in to the calculator wrong for jenny

OpenStudy (anonymous):

you wrote the exponent as 7 but i assume that was a typo and you meant 72 it is easy to make a mistake when calculating this not sure what calculator you use

OpenStudy (anonymous):

Okay so other than that it's right? THANK YOU!! :D :D :D

OpenStudy (anonymous):

yes, take a look at the links i sent wolfram is real easy to use just type it in (i cut and paste what you wrote) and you get the answer

OpenStudy (anonymous):

yw